# Light of Normals: Unified Feature Representation for Universal Photometric Stereo

### v2

@@ Abstract
Universal photometric stereo (PS) aims to recover high-quality surface normals from objects under arbitrary lighting conditions without relying on specific illumination models. Despite recent advances such as SDM-UniPS and Uni MS-PS, two fundamental challenges persist: 1) the deep coupling between varying illumination and surface normal features, where ambiguity in observed intensity makes it difficult to determine whether brightness variations stem from lighting changes or surface orientation; and 2) the preservation of high-frequency geometric details in complex surfaces, where intricate geometries create self-shadowing, inter-reflections, and subtle normal variations that conventional feature processing operations struggle to capture accurately. To address these challenges, we propose LINO-UniPS, featuring three key innovations: (1) learnable light register tokens with global cross-image attention that effectively decouple lighting from normal features; (2) wavelet transform-based sampling that preserves high-frequency surface details; and (3) a normal-gradient confidence loss that prioritizes accurate reconstruction in regions with high normal variations. Additionally, we introduce PS-Verse, a synthetic dataset with graduated geometric complexity and diverse lighting conditions to progressively train models on increasingly challenging scenarios. Extensive experiments demonstrate that our method achieves state-of-the-art performance on public benchmarks while exhibiting strong generalization capabilities across various material properties and lighting scenarios.
Figure 1: (Left) Given multiple images from the same viewpoint under varying lighting, our method outperforms existing universal photometric stereo methods in accuracy and detail, even exceeding the professional 3D scanner. (Right) On the DiliGenT benchmark, with similar decoders, higher feature consistency (SSIM and CSIM) in encoder features corresponds to better accuracy, which is measured by the percentage of pixels with angular error under $9.25^{\circ}$.
@@ 1 Introduction
Photometric stereo (PS) [55] aims to recover surface normals from multiple images under varying lighting conditions. Traditional methods  [29, 30, 24, 40, 19] rely on assumptions of specific light sources or physical models, making them difficult to adapt to complex natural scenes.
Until recently, to overcome the dependence of traditional methods on physical lighting models, the Universal PS approaches [27, 31, 20] employ an encoder-decoder architecture. The encoder extracts high-dimensional representations of pixel-wise, spatially varying illumination from multiple images taken under different lighting conditions, thereby producing a so-called global lighting context that replaces traditional physical parameter modeling. Subsequently, the decoder, functioning as a calibration network [26], leverages this context to estimate surface normals, achieving robust normal recovery in complex and spatially non-uniform lighting environments, thereby significantly advancing the field.
From the design perspective of previous Universal PS methods [31, 27], the features extracted by the encoder are expected to vary significantly with changes in illumination distribution, and the encoder typically processes a larger capacity than the decoder. However, we observe that as the similarity of the lighting context features generated by different encoders increases, the performance of normal estimation also correspondingly improves, as shown in Fig. 1. This indicates that the encoder effectively couples lighting and normal features within the global lighting context, while the decoder is tasked with disentangling the invariant normal features from this varying context, which imposes an increased learning burden on the encoder. Therefore, the encoder should possess the capability to attenuate the coupling between illumination variations and intrinsic surface normal attributes, extracting ideal normal features that are consistent across images and rich in detail from diverse lighting contexts to ensure robust and high-quality normal recovery. The synergy of these two factors ultimately determines the effectiveness of Universal PS in estimating surface normal under complex illumination.
However, Universal PS methods still face significant challenges in encoding such ideal normal features. First, most current methods rely on single image frames and patch-position-aligned attention mechanisms, hindering the full integration of global lighting information from multiple images, resulting in instability and inconsistency in normal feature estimation. Second, commonly used upsampling and downsampling operations (such as bilinear interpolation [28] and pixel shuffling [31]), while computationally efficient, tend to degrade high-frequency details, compromising the fine recovery of normals. Finally, although existing synthetic training datasets cover various types of light sources, they fail to adequately simulate the impact of complex geometry on the spatial distribution of illumination, limiting the model’s generalization ability.
To address the above challenges, we propose LINO-UniPS, a general photometric stereo method based on the ViT [42, 13] architecture. Our approach specifically targets the two fundamental challenges in Universal PS: the deep coupling between illumination and normal features, and the preservation of high-frequency geometric details. Inspired by [9, 53], we use learnable light register tokens to store global illumination (see Fig. 3), effectively alleviating the coupling between lighting and normal features. Combined with a global cross-image attention mechanism, the encoder can generate more consistent and discriminative normal features. By reducing ambiguity in the observed intensity, the model can more effectively distinguish between variations in lighting and actual changes in surface orientation. Furthermore, we use wavelet transform [14, 10] and integrate normal gradient perception loss along with enhanced feature resolution to ensure rich detail in the normal features, thereby improving the recovery of high-quality surface normals.This ensures the preservation of high-frequency components, which are essential for capturing intricate geometries. Moreover, we propose a dataset graded by surface complexity, significantly improving the model’s generalization ability to real-world complex lighting scenarios. Extensive experiments on synthetic and real datasets demonstrate that our method outperforms state-of-the-art techniques in both accuracy and robustness. Overall, our contributions can be summarized as the following three points:
@@
    We use learnable light register tokens for flexible decoupling of illumination and normal features, combined with a global cross-image attention mechanism to enhance multi-view lighting representation and normal feature consistency.
@@
    We adopt wavelet transform for up/downsampling, combined with normal-gradient perception loss and a point-wise decoder, significantly improving fine normal detail recovery.
@@
    We create a synthetic dataset with graded normal complexity to enrich training, and experiments confirm our method’s superior performance and generalization.
@@ 2 Related Work
Calibrated Photometric Stereo: Calibrated PS [55] deduces surface normals by assuming meticulously pre-calibrated illumination, often relying on precise light source parameters. This stringent requirement Initial methods successfully addressed Lambertian surfaces under such known lighting [55]. Subsequent research extended capabilities to handle complex non-Lambertian reflectance, including varied BRDFs and specular highlights [18, 29, 3, 16, 17], albeit always presupposing perfectly known illumination conditions. More recently, deep learning techniques [47, 4, 33, 32] have significantly enhanced the ability to model intricate materials within this calibrated framework. However, the reliance of Calibrated PS on precise light source parameters severely restricts its applications to controlled laboratory environments, thereby limiting its real-world utility.
Figure 2: Overview of the LiNo-UniPS architecture, featuring a Light-Normal Contextual Encoder, Decoder, and loss computation. The encoder includes light registered wavelet-aware downsampler, an enhanced Light-Normal Contextual Attention Module, a DPT-Based [45] Fusion Module, and a WaveUpSampler, which together encode and fuse pixel and wavelet domain features for normal-consistent outputs. The Decoder is similar to SDM-UniPS [31].
@@
Uncalibrated Photometric Stereo: To alleviate photometric stereo’s dependence on pre-calibrated light sources, the task of Uncalibrated Photometric Stereo (Uncalibrated PS) was introduced. Uncalibrated PS methods [21, 2, 38, 6, 50, 34, 36] commonly adopt a two-stage pipeline: an initial stage where a network module estimates the unknown lighting parameters, followed by a second stage where these estimates are utilized as known inputs, like conventional Calibrated PS framework. While such Uncalibrated PS techniques have demonstrated success in handling more complex scenarios [36], their application under unconstrained natural or ambient illumination, encounters significant challenges. These challenges primarily stem from the inherent difficulty in accurately modeling the physics of such arbitrary and often intricate lighting environments.
@@
Universal Photometric Stereo: The Universal PS task, as recently formulated by Ikehata in [28], leverages a purely data-driven approach to solve the photometric stereo problem. This methodology enables operation under arbitrary and unknown lighting environments, critically obviating the need for complex predefined assumptions regarding the illumination. Building on this foundation, methods such as SDM-UniPS [31] further advanced Universal PS by employing an encoder to extract features from multi-illumination images; these features are then fused to create a global lighting context, which a subsequent decoder utilizes in conjunction with pixel-level information to estimate surface normals  [20] has explored multi-scale strategies to address the Universal PS problem; however, such approaches often fall short of reconstructing surface normals that are simultaneously detailed and accurate. We posit that the essence of PS fundamentally involves resolving consistent surface normals from multiple images captured under diverse lighting conditions. Our analysis of prior methods like UniPS and SDM-UniPS indicates that an encoder’s ability to extract features with greater consistency significantly facilitates the decoder’s task of producing higher-fidelity normal maps. Motivated by this observation, we propose our method, LiNO-UniPS, which is designed to learn a superior Unified Feature Representation for Universal Photometric Stereo.
@@ 3 Method
The core task of Universal PS is to accurately recover the surface normal map $N\in\mathbb{R}^{H\times W\times 3}$ of an object or scene from a set of $F$ images $\{I_{f}\}_{f=1}^{F}$, where $I_{f}\in\mathbb{R}^{H\times W\times 3}$, captured under a single viewpoint but multiple unknown lighting conditions [28].
This task faces two key challenges: unknown and spatially varying lighting conditions that are deeply entangled with normal features and difficult to disentangle; and non-Lambertian reflectance from complex surface geometry and materials, making accurate perception difficult. We address these challenges from both methodological and data perspectives. On the methodological side, we propose LINO-UniPS (see Sec. 3.1), whose architecture is shown in Fig. 2. It introduces light registers and a global attention mechanism to decouple lighting from normal features during the encoder stage, as well as wavelet transform and normal gradient-aware loss to enhance the model’s sensitivity to high-frequency details. On the data side, we present the large-scale graded PS-Verse dataset (see Sec. 3.2), which for the first time introduces normal maps to enhance geometric complexity and lighting variation, thereby enriching the diversity of training data.
@@ 3.1 LINO-UniPS
Network Architecture Overview: Given the multi-light images as input to the Light-Normal Contextual Encoder (illustrated in the top pipeline of Fig. 2), they first undergo Light Registered Wavelet-aware DownSampler (leftmost block in the encoder in Fig. 2), which simultaneously generates bilinearly downsampled global components as well as low-frequency and high-frequency wavelet decompositions [10, 14]. Subsequently, the global components and wavelet components are independently converted into patch-based token sequences. These token sequences are prepended with additional light register tokens to aggregate lighting information before being fed into a ViT-based backbone, which is initialized with DINOv2 [42, 13], to extract independent visual representations for each image. Next, the visual representations are processed by four cascaded interleaved attention blocks(central block in the encoder in Fig. 2). Feature tokens, excluding the light register tokens, are then subjected to multi-level aggregation(block before WaveUpSampler in Fig. 2). Using WaveUpSampler(rightmost block in Fig. 2), features from the wavelet domain and pixel domain are fused to produce what we term the Light-Normal Contextual representation. Finally, this Light-Normal Contextual representation is decoded by a decoder to predict the surface normal map. To maintain clarity, we emphasize the discussion of high-level concepts instead of detailed explanations. A full description of the detailed implementation can be found in the appendix.
Figure 3: We visualize the attention maps of lighting registers on the encoder’s final-layer feature maps for both real (left) and synthetic (right) data. The first and second rows correspond to HDRI and point light registers, respectively, showing attention regions that rotate from left to right, demonstrating the model’s ability to capture light from different directions. The HDRI register exhibits broader attention regions, while the point light register focuses more sharply, consistent with the low-frequency nature of HDRI and the high-frequency characteristics of point lights.
@@
Figure 4: Features from different method encoders; rightmost column is variance.
@@
Figure 5: Results of object inference with masks in Luces and DiLiGenT datasets.
Strengthening Lighting-Normal Decoupling: Previous universal photometric stereo methods [28, 31] encode spatially varying global lighting context per pixel from multi-illuminated images and pass it as physical lighting parameters to a self-calibrating decoder for normal prediction. Usually, the encoder is far more powerful than the decoder. However, as observed in Fig. 5, this approach can lead to the encoder deeply coupling lighting and normal features, putting too much pressure on the decoder to disentangle them. This motivates our approach to decouple lighting and normal features within the encoder, allowing the decoder to focus solely on estimating normals from consistent normal features, thus easing the learning process.The success of this decoupling is quantitatively shown by improved CSIM and SSIM scores in Tab. 2 and Tab. 3.
Specifically, our training process consists of three key components: Light Registers Tokens, Interleaved Attention, and the Light Aligner. Inspired by the register mechanism in DINOv2 [42], we designed Light Registers to collect global information during model training, which can be discarded during inference. Through supervised learning, the encoder’s output is guided to encode features consistent with surface normals, achieving an initial decoupling of lighting and normal features. In contrast, UniPS and SDM-UniPS only use Frame Attention and Light Axis Attention to enable local information flow and global lighting context communication, which is insufficiently efficient to disentangle the complex coupling of lighting variations and normal features. Inspired by VGGT [53], we introduce a global attention mechanism that attends to all input image tokens, facilitating efficient interaction between global lighting context and normal features, thus achieving more effective information decoupling. Additionally, drawing on the feature alignment methods used in VAVAE [59] and REPA [60] to accelerate generative model training, we encode three of the light registers along with information about the types of light sources in the training dataset—including the HDRI map, as well as the intensities, sizes, and positions of area lights and point lights (see Fig. 2)—and constrain each of them separately with a cosine similarity loss. We summarize the form of the loss functions as follows:
$\mathcal{L}_{\text{light}}=\lambda_{1}\mathcal{L}_{\text{hdri}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{area}}$
Where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters. Through the combined effect of these three components, we significantly enhance the decoupling of lighting and normal features; detailed results are shown in Fig. 5 and Tab. 3.
Extracting Fine-grained Contexts: Current feature extraction techniques often discard high-frequency surface details, which are vital for accurate normal reconstruction. Especially in complex surface regions, the precision and detail of predicted surface normals improve significantly when the encoder can capture finer-grained contextual information. Although prior methods have recognized this challenge, they have not effectively resolved it. UniPS [28] extracts features from downsampled images, inherently introducing blur. While SDM-UniPS [31] attempted to mitigate this blur through a "split-and-merge" operation, it disrupted the original high-frequency semantics [41]. To address these shortcomings and overcome these limitations, enabling our encoder to effectively capture and preserve fine-grained details, we introduce the wavelet transform and a normal gradient perception loss.
To enable the encoder to extract finer-grained contexts, we replace the original downsampling operation with discrete wavelet transform [10] to extract the high-frequency and low-frequency components of multi-light images, thereby mitigating information loss during downsampling. At the same time, to retain global image-domain semantic information, similar to SDM-UniPS, we also preserve bilinear downsampling. Refer to Fig. 2 for the light-registered wavelet-aware downsampler. Additionally, during feature upsampling, we use discrete wavelet inverse transform in the wavelet upsampler to upsample features in the wavelet domain.
@@
Table 1: Comparison of PS-Verse with Other Photometric Stereo Datasets. The term ‘Normal‘ indicates whether the normal map is used to enhance surface details during data rendering. ‘Light‘ denotes the presence of light type annotations, while ‘Complexity‘ quantifies the average magnitude of surface normal gradients within the dataset, where higher values signify greater geometric complexity.
Dataset Intrinsic Normal Light Complexity # Shapes # HDRI # Scenes
CyclesPS-Train [25] 4.9 15 0 45
PS-Wild [28] 3.5 410 31 10,099
PS-Mix [31] 11.5 410 31 34,927
PS-Uni MS-PS [20] 8.6 11,000 1100 100,000
@@
To overcome the limitations of existing methods in handling complex geometric and richly textured regions, we introduce a Normal Gradient Perception Loss. This loss guides the network by using the predicted normal gradient to generate a confidence map that weights the normal prediction error, thereby emphasizing high-frequency detail areas. The combined loss is formulated as:
$\mathcal{L}_{N}=\lambda_{4}\sum(N-\tilde{N})^{2}\odot C+\lambda_{5}\sum(\tilde{G}-G)^{2},$
where the confidence map $C=e^{\tilde{G}}$ is derived from the predicted normal gradient $\tilde{G}$, $G=\nabla N$ is the ground truth gradient, and $\lambda_{4},\lambda_{5}$ are weighting coefficients balancing the normal prediction and gradient supervision terms. This design effectively enhances the network’s sensitivity to fine surface details in challenging regions.

### v4

@@ Abstract
Universal photometric stereo (PS) is defined by two factors: it must (i) operate under arbitrary, unknown lighting conditions and (ii) avoid reliance on specific illumination models. Despite progress (e.g., SDM UniPS), two challenges remain. First, current encoders cannot guarantee that illumination and normal information are decoupled. To enforce decoupling, we introduce LINO UniPS with two key components: (i) Light Register Tokens with light alignment supervision to aggregate point, direction, and environment lights; (ii) Interleaved Attention Block featuring global cross-image attention that takes all lighting conditions together so the encoder can factor out lighting while retaining normal-related evidence. Second, high-frequency geometric details are easily lost. We address this with (i) a Wavelet-based Dual-branch Architecture and (ii) a Normal-gradient Perception Loss. These techniques yield a unified feature space in which lighting is explicitly represented by register tokens, while normal details are preserved via wavelet branch. We further introduce PS-Verse, a large-scale synthetic dataset graded by geometric complexity and lighting diversity, and adopt curriculum training from simple to complex scenes. Extensive experiments show new state-of-the-art results on public benchmarks (e.g., DiLiGenT, Luces), stronger generalization to real materials, and improved efficiency; ablations confirm that Light Register Tokens + Interleaved Attention Block drive better feature decoupling, while Wavelet-based Dual-branch Architecture + Normal-gradient Perception Loss recover finer details.
Figure 1: (left) Given multi-light images from a fixed viewpoint, LINO UniPS recovers sharper, more faithful normals than UniPS/SDM-UniPS and visually rivals a 3D scanner. (right) On the DiLiGenT, a clear correlation exists between the consistency of encoder features (CSIM/SSIM) and the final reconstruction accuracy (1/MAE).
@@ 1 Introduction
Photometric stereo (PS) Woodham (1980) aims to recover surface normals from multiple images under varying lighting conditions. Traditional methods Ikehata et al. (2012; 2014); Ikehata (2018); Mo et al. (2018); Haefner et al. (2019) rely on assumptions of specific light sources or physical models, making them difficult to adapt to complex natural scenes.
Recent Universal PS methods Ikehata (2022; 2023); Hardy et al. (2024) have alleviated this reliance by replacing explicit physical modeling with a learnable encoder-decoder architecture. The encoder extracts representations of pixel-wise, spatially varying illumination from multiple input images and produces a global lighting context, while the decoder, acting as a calibration network Ikehata (2021), transforms this context into normal predictions. This paradigm achieves robust normal estimation in diverse and spatially non-uniform lighting environments, substantially advancing the field. Yet, achieving reliable performance in this setting is more challenging than it seems.
The first challenge lies in the ineffective decoupling of illumination and normal cues. Existing encoders jointly process lighting and normal, but without explicit illumination representation the decoder inherits unstable features, producing inconsistent normal predictions. Interestingly, we observe that when encoders produce more similar lighting features across different inputs, the accuracy of normal estimation improves (Fig. 1 right), suggesting that explicitly factoring out illumination is crucial for normal recovery. The second challenge is the loss of geometry details. Conventional up/downsampling operations (e.g., bilinear interpolation Ikehata (2022), pixel shuffle Ikehata (2023)) either smooth fine details or disrupt high-frequency semantics Odena et al. (2016), leading to degraded normal quality, especially in regions with complex variation (see Fig. 1 left).
In this paper, we propose LINO UniPS, a ViT-based framework that explicitly decouples illumination from surface cues while preserving geometric detail. First, we introduce learnable Light Register Tokens to capture illumination information. To handle diverse lights, we design three distinct types of tokens, each dedicated to a specific light source: point, direction light, and environment (Env) lights. Our key innovation is an explicit light alignment supervision, which guides each register token to learn the characteristics of its corresponding illumination type. Second, we introduce an Interleaved Attention Block to process the image features alongside these specialized registers. This block employs a global cross-image attention mechanism that simultaneously attends to all tokens from all lighting conditions. The synergy between the Light Register Tokens and the Interleaved Attention Block enables the decoupling of lighting and normal cues, allowing our encoder to yield a unified feature representation. In parallel, we introduce two components to ensure the preservation of fine-scale geometry: a Wavelet-based Dual-branch Architecture, which introduces the wavelet domain to help the model better attend to high-frequency details, and a Normal-gradient Perception Loss that more heavily penalizes errors in high-frequency regions during training, thereby further refining local predictions. We further construct PS-Verse, a new dataset graded by surface complexity and lighting diversity, which enhances robustness and generalization under challenging real-world lighting. Together, these designs enable state-of-the-art accuracy, robustness, and generalization across synthetic and real benchmarks.
Overall, our contributions can be summarized as the following three points:
@@
    We introduce Light Register Tokens and an Interleaved Attention Block to explicitly decouple illumination from normal features, thereby yielding a unified feature representation.
@@
    We adopt a Wavelet-based Dual-branch Architecture and a Normal-gradient Perception Loss, which together substantially improve the reconstruction of fine-grained geometric details.
@@
    We build PS-Verse, a synthetic dataset graded by surface complexity and lighting diversity, and demonstrate superior accuracy and generalization through curriculum training.
@@ 2 Related Work
Calibrated Photometric Stereo: Calibrated PS Woodham (1980) deduces surface normals by assuming meticulously pre-calibrated illumination, often relying on precise light source parameters. This stringent requirement limits the applicability of early methods. Initial methods successfully addressed Lambertian surfaces under such known lighting Woodham (1980). Subsequent research extended capabilities to handle complex non-Lambertian reflectance, including varied BRDFs and specular highlights Grossberg & Nayar (2003); Chandraker et al. (2005); Goldman et al. (2005; 2010); Ikehata et al. (2012), albeit always presupposing perfectly known illumination conditions. More recently, deep learning techniques Jung et al. (2015); Santo et al. (2017); Chen et al. (2018); Ju et al. (2021) have significantly enhanced the ability to model intricate materials within this calibrated framework. However, the reliance of Calibrated PS on precise light source parameters severely restricts its applications to controlled laboratory environments, thereby limiting its real-world utility.
@@
Uncalibrated Photometric Stereo: To alleviate photometric stereo’s dependence on pre-calibrated light sources, the task of Uncalibrated Photometric Stereo (Uncalibrated PS) was introduced. Uncalibrated PS methods Hayakawa (1994); Belhumeur & Kriegman (1996); Mallick et al. (2005); Chen et al. (2020); Taniai & Maehara (2018); Kaya et al. (2021); Lichy et al. (2022) commonly adopt a two-stage pipeline: an initial stage where a network module estimates the unknown lighting parameters, followed by a second stage where these estimates are utilized as known inputs, like conventional Calibrated PS frameworks. While such Uncalibrated PS techniques have demonstrated success in handling more complex scenarios Lichy et al. (2022), their application under unconstrained natural or ambient illumination encounters significant challenges. These challenges primarily stem from the inherent difficulty in accurately modeling the physics of such arbitrary and often intricate lighting environments.
@@
Universal Photometric Stereo: The Universal PS task, as recently formulated by Ikehata in Ikehata (2022), leverages a purely data-driven approach to solve the photometric stereo problem. This task enables operation under arbitrary and unknown lighting environments, critically obviating the need for complex predefined assumptions regarding the illumination. Building on this foundation, methods such as SDM UniPS Ikehata (2023) further advanced Universal PS by employing an encoder to extract features from multi-illumination images; these features are then fused to create a global lighting context, which a subsequent decoder utilizes in conjunction with pixel-level information to estimate surface normals. Uni MS-PS Hardy et al. (2024) has explored multi-scale strategies to address the Universal PS problem; however, such approaches often fall short of reconstructing surface normals that are simultaneously detailed and accurate. Our analysis of prior methods like UniPS and SDM UniPS indicates that an encoder’s ability to extract features with greater consistency significantly facilitates the decoder’s task of producing higher-fidelity normal maps. Motivated by this observation, we propose our method, LINO UniPS, which is designed to learn a superior Unified Feature Representation for Universal Photometric Stereo. Furthermore, we adopt a Wavelet-based Dual-branch Architecture and a Normal-gradient Perception Loss to address the challenge of detailed reconstruction in high-frequency regions.
@@ 3 Method
The core task of Universal PS is to accurately recover the surface normal map $N\in\mathbb{R}^{H\times W\times 3}$ of an object or scene from a set of $F$ images $\{I_{f}\}_{f=1}^{F}$, where $I_{f}\in\mathbb{R}^{H\times W\times 3}$, captured under a single viewpoint but multiple unknown lighting conditions Ikehata (2022).
As illustrated in Fig. 2, our method processes multi-light images through an encoder-decoder architecture. Within the encoder, the pipeline consists of three main stages: 1) First, the input images undergo processing by the Wavelet Feature Extractor. This module simultaneously generates downsample components and low- and high-frequency wavelet components via wavelet decomposition Daubechies (1990); Finder et al. (2024), and then converts both components into independent token sequences through the ViT-based backbone Dosovitskiy et al. (2021). 2) These sequences then enter the Light Registered Attention Module, where they are prepended with specialized Light Register Tokens (Point, Direction, Env) to aggregate global illumination, before being processed by cascaded Interleaved Attention Blocks. 3) Next, these tokens, excluding the Light Register Tokens, are fed into the Wavelet Aggregator. Here, features from the downsample and wavelet branches are fused to produce the unified feature representation, which is finally passed to a decoder for normal prediction.
The training is guided by two supervisory signals. First, a light alignment loss ensures that the register tokens capture the characteristics of their corresponding illumination sources; this loss is assigned a small weight, allowing it to serve as a regularizer. Second, a Normal Gradient Perception Loss enhances the model’s ability to reconstruct fine-grained geometric details. Furthermore, this section only provides a high-level overview; please refer to appendix A1.1 for details.
Figure 2: Overview of the LINO UniPS architecture. The encoder includes a Wavelet Feature Extractor, a Light Registered Attention Module and a Wavelet Aggregator, which together encode and fuse wavelet and downsample domain features for obtaining unified feature. The decoder is similar to SDM UniPS Ikehata (2023).
@@ 3.1 LINO-UniPS
Despite recent progress in this field, prior methods still suffer from two primary challenges.
Challenge 1: The encoder fails to effectively decouple illumination and normal information.
This failure forces the decoder to perform this disentanglement in addition to its primary prediction task. This creates a paradoxical workflow, as the more powerful encoder partially offloads the most difficult disentanglement to the weaker decoder. Motivated by this, our approach empowers the encoder to generate consistent, disentangled normal features. This simplifies the decoder’s role to that of a simple refiner, easing the learning process. To achieve this, our method introduces two core components for this purpose: Light Register Tokens and Interleaved Attention Block.
Light Register Tokens: Inspired by the register mechanism in DINO Darcet et al. (2024), we introduce Light Register Tokens to aggregate global illumination information. To account for varied illumination types, we design distinct register tokens for point, direction, and environment (Env) lights. Our key innovation, and a crucial departure from DINO’s unsupervised registers, is the introduction of an explicit light alignment supervision.
Drawing on the feature alignment methods used in VAVAE Yao et al. (2025) and REPA Yu et al. (2025) to accelerate generative model training, we encode the three types of Light Register Tokens to align with information about the light sources in the training dataset, including the point, direction and environment lights, (see Fig. 2), and supervise each of them separately with a cosine similarity loss. We summarize the form of the loss functions as follows:
$\mathcal{L}_{\text{light}}=\lambda_{1}\mathcal{L}_{\text{env}}+\lambda_{2}\mathcal{L}_{\text{point}}+\lambda_{3}\mathcal{L}_{\text{direction}}$
where $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are hyperparameters.The settings for these hyperparameters and the detailed formulations of the $\mathcal{L}_{\text{env}}$, $\mathcal{L}_{\text{point}}$, and $\mathcal{L}_{\text{direction}}$ are provided in appendix A1.1.7 and appendix A1.1.3.
Through light alignment supervision, these tokens learn to capture and encode their respective illumination information, thereby achieving decoupling of lighting from normal features. This is demonstrated by the phenomena observed in Fig. 3: the register tokens show specialized focus: Direction and Env register tokens attend to broad regions, while the Point register tokens focuse more sharply. This aligns with the expected low- and high-frequency characteristics of these light types, respectively, confirming our registers’ effectiveness.
Interleaved Attention Block: Previous methods like UniPS and SDM UniPS rely solely on Frame Attention (Frame) Dosovitskiy et al. (2021) and Light Axis Attention (Light) Ikehata (2023). While these mechanisms facilitate local information flow, they are insufficient for effectively decoupling lighting variations and normal features. Inspired by VGGT Wang et al. (2025), our Interleaved Attention block introduces a more powerful global cross-image attention mechanism (Global) that attends to all input image tokens simultaneously. Specifically, our block applies four attention layers in an interleaved sequence: Frame $\rightarrow$ Light $\rightarrow$ Global $\rightarrow$ Light (see Fig.2).
The key advantage of this design is its ability to aggregate and fuse information across multiple hierarchical levels. Light-axis attention operates at the patch level, frame attention captures inter-image context, and our global attention integrates information at the cross-image level. This multi-level awareness enables the model to build a holistic understanding of the global illumination from local to global scales, thereby achieving a better decoupling from the intrinsic normal features.
Challenge 2: The model fails to preserve high-frequency geometric details.
Accurate normal reconstruction is highly dependent on preserving high-frequency geometric details, yet current feature extraction techniques often discard them. Although prior methods recognize this issue, their solutions are inadequate. UniPS Ikehata (2022) extracts features from downsampled images, which inherently introduces blurring. While SDM UniPS Ikehata (2023) attempts to mitigate this with a "split-and-merge" operation, this process can disrupt high-frequency semantics Odena et al. (2016). To overcome these limitations, we propose a two-pronged approach: a Wavelet-based Dual-branch Architecture to preserve information during feature extraction, and a Normal Gradient Perception Loss to guide the model’s focus towards these fine details during training.
Wavelet-based Dual-branch Architecture: To mitigate information loss during downsampling, our dual-branch uses the discrete wavelet transform Daubechies (1990). This allows us to decompose multi-light images into their high- and low-frequency components. At the same time, to retain global image-domain semantic information, we maintain a parallel branch that performs downsampling. During feature upsampling, an inverse wavelet transform reconstructs features from the wavelet domain, ensuring detail preservation throughout the network.
Figure 3: The attention maps for our different Light Register Tokens, derived from the encoder’s final-layer feature maps for both real (left) and synthetic (right) data.
Normal Gradient Perception Loss: To further guide the network’s focus towards complex geometric and richly textured regions, we introduce a Normal Gradient Perception Loss ($\mathcal{L}_{\text{n}}$). Instead of treating all pixels equally, this loss uses the predicted normal gradient ($\tilde{G}$) to generate a confidence map ($C=e^{\tilde{G}}$) that amplifies the error signal in high-frequency areas. The loss is a weighted sum of this confidence-weighted reconstruction error and a gradient supervision term:
$\mathcal{L}_{n}=\lambda_{4}\sum(N-\tilde{N})^{2}\odot C+\lambda_{5}\sum(\tilde{G}-G)^{2},$ (1)
Here, the first term penalizes the difference between the predicted normal ($\tilde{N}$) and ground truth ($N$), weighted by the confidence map $C$. The second term directly supervises the predicted gradient ($\tilde{G}$) against the ground truth gradient ($G=\nabla N$). The coefficients $\lambda_{4}$ and $\lambda_{5}$ balance these two objectives. This design makes the network explicitly sensitive to fine surface details, significantly improving reconstruction quality in challenging regions.
@@
Figure 4: Features from different methods’ encoders; rightmost column is variance.
@@
Figure 5: Results of object inference with masks in Luces and DiLiGenT. Using 16 input images.
@@
Table 1: Comparison of PS-Verse with Other Photometric Stereo Datasets. The terms ‘Intrinsic‘, ‘Normal‘, and ‘Light‘ indicate the availability of material parameters (albedo, etc.), the use of normal maps for rendering detail, and the presence of light type annotations. ‘Complexity‘ quantifies the average magnitude of surface normal gradients, where higher values denote greater geometric intricacy. ‘Shapes‘, ‘Env‘, and ‘Scenes‘ are the respective counts of 3D models, HDRI environment maps, and rendered scenes.
Dataset Intrinsic Normal Light Complexity # Shapes # Env # Scenes
CyclesPS-Train Ikehata (2018) 4.9 15 0 45
PS-Wild Ikehata (2022) 3.5 410 31 10,099
PS-Mix Ikehata (2023) 11.5 410 31 34,927
PS-Uni MS-PS Hardy et al. (2024) 8.6 11,000 1100 100,000